OpenStudy (anonymous):

Number Thoery by Geogre Andrews page 68. Explain the part where it says "Let ni = M/mi. Since no two of the mi have a common factor, we see that gcd(ni, mi) = 1" Why?

OpenStudy (mathstudent55):

Wasn't that the supposition at the statement of the theorem above?

OpenStudy (anonymous):

uhm.. no. The assumption was no two of the mi have a common factor. M = m1 * m2 * ...* ms ni = M / mi, the claim is gcd(ni,mi) = 1

jimthompson5910 (jim_thompson5910):

It has to do with how M is set up M = m1*m2*m3 ... *ms and how the mi's are pairwise coprime eg: If i = 2, then ni = M/mi n2 = M/m2 n2 = m1*m3*m4*...*ms n2 and m2 have no factors in common other than 1 because m2 and m1*m3*m4*...*ms have nothing in common other than 1

OpenStudy (anonymous):

@jim_thompson5910 oh I see now. It has something to do with finding the greatest common factor of two numbers by prime factorization. (something I forgot I learned in grades school lol)

jimthompson5910 (jim_thompson5910):

that is correct sometimes you'll see `gcd(a,b)` to mean `the greatest common factor of a and b` other times you may see `(a,b)` to mean the same thing

OpenStudy (anonymous):

@jim_thompson5910 awesome. Thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome